### v1

# Collaborative Deterministic-Diffusion Model for Probabilistic Urban Spatiotemporal Prediction

Abstract.
Accurate prediction of urban spatiotemporal dynamics is essential for enhancing urban management and decision-making. Existing spatiotemporal prediction models are predominantly deterministic, focusing on primary spatiotemporal patterns. However, those dynamics are highly complex, exhibiting multi-modal distributions that are challenging for deterministic models to capture. In this paper, we highlight the critical role of probabilistic prediction in capturing the uncertainties and complexities inherent in spatiotemporal data. While mainstream probabilistic models can capture uncertainty, they struggle with accurately learning primary patterns and often suffer from computational inefficiency. To address these challenges, we propose CoST, which collaborates deterministic and probabilistic models to improve both predictive accuracy and the ability to handle uncertainty. To achieve this, we design a mean-residual decomposition framework, where the mean value is modeled by a deterministic model, and the residual variations are learned by a probabilistic model, specifically diffusion models. Moreover, we introduce a scale-aware diffusion process, which better accounts for spatially heterogeneous dynamics across different regions. Extensive experiments on eight real-world datasets demonstrate that CoST significantly outperforms existing methods in both deterministic and probabilistic metrics, achieving a 20% improvement with low computational cost. CoST bridges the gap between deterministic precision and probabilistic uncertainty, making a significant advancement in the field of urban spatiotemporal prediction.
Probabilistic prediction, deterministic models, diffusion models
1. Introduction
Figure 1. Comparison of existing models with our mean-residual decomposition framework.
Accurate prediction of urban spatiotemporal dynamics is crucial for urban management, efficiency optimization, and the development of smart cities (Zhang et al., 2017; Xie et al., 2020; Jin et al., 2023a; Yuan et al., 2024b). These urban spatiotemporal data encapsulate complex patterns while also exhibiting randomness (Sheng et al., 2025; Chirigati et al., 2016; Li et al., 2020; Behnisch and Ultsch, 2009; Zhang et al., 2017). Taking traffic flow data as an example, it displays clear temporal and spatial patterns: it peaks during rush hours and is influenced by neighboring regions, while also being prone to random fluctuations caused by events like accidents or extreme weather.
Mainstream urban spatiotemporal prediction methods mainly rely on deterministic models (Shao et al., 2022; Yan et al., 2018; Yuan et al., 2024a, b; Zhang et al., 2017). These models typically use mean absolute error ($\mathcal{L}_{1}$) or squared error ($\mathcal{L}_{2}$) between predicted and true values as the loss function. This allows to learn the expected value and capture primary spatiotemporal patterns. However, the data often follow a multi-modal distribution (Gao et al., 2024; Han et al., 2022), where multiple states can coexist at a given spatiotemporal point. For example, in a scenario where a road segment with frequent accidents has an accident rate $p_{1}$, with peak traffic volumes $y_{1}$ during normal hours and $y_{2}$ during accidents, deterministic models learn a conditional mean: $\mathbb{E}[y|x]=p_{1}\times y_{2}+(1-p_{1})\times y_{1}$. However, this underestimates the resources needed in congested areas and over-allocates them to free-flowing ones. In contrast, probabilistic methods can address this by modeling multi-modal distributions.
Existing probabilistic models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Gao et al., 2022), Variational Autoencoders (VAEs) (Kingma, 2013; Li et al., 2022), and Diffusion Models (Ho et al., 2020; Tashiro et al., 2021), have shown promising performance. These methods aim to learn the full data distribution, capturing both primary patterns (e.g., long-term temporal trends and global spatial patterns) and non-linear variations (e.g., random fluctuations caused by incidental events). However, their focus on modeling uncertainty often limits their ability to accurately capture the primary patterns. Additionally, these models often suffer from inefficiency in real-world deployment. For example, GANs struggle with training challenges and mode collapse (Goodfellow et al., 2020; Brophy et al., 2021), while diffusion models require multiple denoising steps for sampling, increasing computational costs (Tashiro et al., 2021; Rühling Cachay et al., 2023; Yuan et al., 2023).
Therefore, we aim to leverage the strengths of both deterministic and probabilistic models to enhance predictive accuracy while effectively capturing uncertainty. However, this is not trivial and faces three challenges: (i) Spatiotemporal data not only exhibits primary spatiotemporal patterns but also involves random fluctuations. (ii) The multi-modal distributions across different spatiotemporal locations are inherently heterogeneous. (iii) Large-scale urban computing and decision-making applications require both computationally efficient and scalable prediction models.
To address the aforementioned challenges, we propose CoST that Collaborates the deterministic model and probabilistic model for urban SpatioTemporal prediction. As illustrated in Figure 1, we leverage an advanced deterministic spatiotemporal prediction model to capture the primary patterns and predict the conditional mean $\mathbb{E}(y|x)$. On this basis, we employ the diffusion model to model the residual distribution $p(r|x)=p(y-\mathbb{E}(y|x)|x)$, compensating for the limitations of the mean prediction. Additionally, we propose a scale-aware diffusion process to handle the spatial heterogeneity. To thoroughly evaluate the performance of probabilistic predictions, we conduct experiments on eight real-world datasets, encompassing crowd flows, cellular network traffic, vehicle flow, and traffic speed. We use both deterministic metrics (MAE, RMSE) and probabilistic metrics (CRPS, QICE, IS) for evaluation. Experimental results demonstrate that our model outperforms existing approaches in both predictive performance and computational efficiency. In summary, our main contributions are as follows:
  • We propose to collaborate deterministic and probabilistic models for probabilistic urban spatiotemporal prediction, leveraging their strengths in both accuracy and uncertainty handling.
  • We propose CoST, a mean-residual decomposition framework where the mean is modeled deterministically and residual variations are learned probabilistically using diffusion models. Additionally, we introduce a scale-aware diffusion process to capture spatial variability in residual distributions.
  • We conduct experiments on 8 real-world datasets using comprehensive metrics for evaluation. Results demonstrate that CoST outperforms existing approaches across multiple key metrics, achieving 20% performance improvement with high efficiency.
2. Related Work
2.1. Urban Spatiotemporal Prediction
Spatiotemporal prediction in urban environments (Zhang et al., 2017; Yuan et al., 2024a; Yu et al., 2017) aims to model and forecast the dynamics of urban data, such as traffic flow, cellular network traffic, and energy consumption. Previously, researchers have employed statistical methods (e.g., ARIMA (Bouznad et al., 2020)) and traditional machine learning models (e.g., SVM (Ohashi and Torgo, 2012), GP (Senanayake et al., 2016), and RF (Hengl et al., 2018)) for prediction. Recently, a variety of neural network-based models have been proposed for this task, which can be categorized into deterministic and probabilistic prediction.
2.2. Deterministic Prediction
Deterministic prediction of spatiotemporal data is a point estimation approach that assumes a deterministic prediction can be made given the same historical observations. These models are typically trained using loss functions such as MSE or MAE, ensuring that the model learns to predict the conditional mean $\mathbb{E}[y|x]$ to capture the primary patterns. Existing deep learning-based deterministic forecasting models include MLP-based (Shao et al., 2022; Qin et al., 2023; Zhang et al., 2023), CNN-based (Li et al., 2017; Liu et al., 2018; Zhang et al., 2017), and RNN-based (Bai et al., 2019b; Lin et al., 2020; Wang et al., 2017, 2018) architectures, which are commonly used for their computational efficiency. Additionally, GNN-based models (Bai et al., 2019a, a, 2020; Geng et al., 2019; Jin et al., 2023a) have been developed to capture spatial dependencies in graph-structured data, while Transformer-based models (Chen et al., 2022, 2021a; Jiang et al., 2023; Yu et al., 2020) excel at capturing complex dynamic patterns.
2.3. Probabilistic Prediction
The core of probabilistic prediction lies in constructing a conditional probability distribution $p(y|x)$, which quantifies the uncertainty of the prediction rather than merely providing a single deterministic value (Yang et al., 2024; Tashiro et al., 2021). Unlike deterministic prediction models, these models assess the reliability of the predictions through probabilistic representations like quantiles or confidence intervals. Early applications typically employ Bayesian networks. In recent years, with the development of generative models, approaches based on GAN (Jin et al., 2022; Saxena and Cao, 2019; Zhang et al., 2021), VAE (Chen et al., 2021b; De Miguel et al., 2022; Zhou et al., 2020), and diffusion models (Tashiro et al., 2021; Chai et al., 2024; Lin et al., 2024) have made significant advancements, owing to their powerful data distribution modeling capabilities. However, deterministic models still dominate in related research, and probabilistic prediction models remain relatively underexplored. This study aims to leverage the strengths of both deterministic and probabilistic models to achieve efficient probabilistic predictions.
3. Preliminary
3.1. Problem Formulation
Urban spatiotemporal prediction aims to predict future variations based on historical observations. The urban spatiotemporal data is typically represented as a multi-dimensional tensor $\mathbf{x}\in\mathbb{R}^{T\times K\times C}$, where $T$ denotes the temporal dimension, $K$ represents the spatial dimension, and $C$ is the feature dimension. The format typically includes two types: grid-based data and graph-based data.
For grid-based data, the spatial dimension $K$ can be expressed in a two-dimensional form as $H\times W$, where $H$ and $W$ denote the height and width, respectively.
For graph-based data, $K$ represents the number of nodes in the graph structure. These nodes are located within a spatial topological structure denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$. Here, $\mathcal{V}$ is the set of all nodes. $\mathcal{E}$ is the set of edges connecting the nodes. In addition, $\mathcal{A}$ is the adjacency matrix of graph $\mathcal{G}$. Its elements $a_{ij}$ show if there’s an edge between node $i$ and $j$ in $\mathcal{V}$, $a_{ij}=1$ when there’s an edge and $a_{ij}=0$ otherwise.
3.2. Diffusion-based Probabilistic Prediction
The diffusion-based urban spatiotemporal model is a conditional probabilistic model, consisting of a forward process and a reverse process. In the forward process, noise is added incrementally to target data $\mathbf{x}^{ta}_{0}$ according to a predefined noise schedule $\{\beta_{n}\}_{t=1}^{N}$, gradually transforming the data distribution into a standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. At any diffusion step, the corrupted target data can be computed using the one-step forward equation:
(1) $\mathbf{x}_{n}^{ta}=\sqrt{\bar{\alpha}_{n}}\mathbf{x}_{0}^{ta}+\sqrt{1-\bar{\alpha}_{n}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
where $\bar{\alpha}_{n}=\prod_{i=1}^{n}\alpha_{i}$ and $\alpha_{n}=1-\beta_{n}$. In the reverse process, we first sample $\mathbf{x}^{ta}_{N}$ from the standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ and denoise it through the following Markov process:
(2) $p_{\theta}(\mathbf{x}_{0:N}^{ta}):=p(\mathbf{x}_{N}^{ta})\prod_{n=1}^{N}p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta},\mathbf{x}_{0}^{co}),$
$p_{\theta}(\mathbf{x}_{n-1}^{ta}|\mathbf{x}_{n}^{ta}):=\mathcal{N}(\mathbf{x}_{n-1}^{ta};\mu_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co}),\Sigma_{\theta}(\mathbf{x}_{n}^{ta},n)),$
$\mu_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co})=\frac{1}{\sqrt{\bar{\alpha}_{n}}}\left(\mathbf{x}_{n}^{ta}-\frac{\beta_{n}}{\sqrt{1-\bar{\alpha}_{n}}}\mathbf{\epsilon}_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co})\right)$
where the variance $\Sigma_{\theta}(\mathbf{x}_{n}^{ta},n)=\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}\beta_{n}$, and $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ is predicted by the denoising network which is trained by the loss function below:
(3) $\mathcal{L}(\theta)=\mathbb{E}_{n,\mathbf{x}_{0},\mathbf{\epsilon}}\left[\left\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{n}^{ta},n|\mathbf{x}_{0}^{co})\right\|_{2}^{2}\right].$
Figure 2. The overview of CoST. (a) illustrates the pretraining process of the deterministic prediction model. (b) demonstrates the computation of the Customized Fluctuation Scale. (c) presents the overall framework of the mean-residual decomposition.
3.3. Evaluation of Probabilistic Prediction
To comprehensively evaluate probabilistic prediction, it is essential to consider two key dimensions:
  • Data Distribution—the predicted distribution should align with the empirical data distribution.
  • Prediction Usability—A reliable prediction interval has a high probability of containing the true future values, while a sharp interval provides a more precise estimate.
Existing studies predominantly rely on Continuous Ranked Probability Score (CRPS) to provide a measure of the distribution discrepancy, while deterministic metrics such as MAE and RMSE to capture point estimation errors. However, these metrics fail to systematically evaluate: (i) Whether the learned distribution accurately covers the data across different quantile intervals, particularly when dealing with multi-modal distributions; (ii) Whether the interval width appropriately reflects real uncertainty. To address these challenges, we introduce two more evaluation metrics, Quantile Interval Coverage Error (QICE) and Interval Score (IS).
QICE
QICE is first proposed in (Han et al., 2022) for regression tasks. This metric is intended to assess the mean absolute deviation between the proportion of true data points within each quantile interval (QI) and the ideal proportion. Specially, for each sample point $y$, after obtaining a sufficient number of predicted values $\hat{y}$, these predictions are divided into $M_{\text{QIs}}$ equally-sized QIs. Then the QICE is computed according to the following formula:
(4) $\text{QICE}:=\frac{1}{M_{\text{QIs}}}\sum_{m=1}^{M_{\text{QIs}}}\left|r_{m}-\frac{1}{M_{\text{QIs}}}\right|,$
$r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathbb{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$
where $\hat{y}_{n}^{\text{low}_{m}}$ and $\hat{y}_{n}^{\text{high}_{m}}$ represents the lower and upper bound of the $m-$th quantile interval for $y_{n}$. When the true distribution aligns with the predicted distribution, $\frac{1}{M_{\text{QIs}}}$ of the observation data will fall into each QI, achieving an optimal QICE value of 0. Smaller values of QICE indicate a closer match between the learned distribution and the true distribution.
IS
The IS is a scoring metric designed to evaluate the quality of probabilistic prediction intervals (PIs). It is defined and systematically studied in (Gneiting and Raftery, 2007). This metric jointly considers two critical aspects: the sharpness and coverage ability of the prediction interval. The computation of IS can be formulated as follows:
(5) $IS:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{\alpha}-l_{n}^{\alpha})+\frac{2}{\alpha}(l_{n}^{\alpha}-y_{n})\mathbb{1}_{y_{n}<l_{n}^{\alpha}}+\frac{2}{\alpha}(y_{n}-u_{n}^{\alpha})\mathbb{1}_{y_{n}>u_{n}^{\alpha}}\right],$
where $u_{n}^{\alpha}$ and $l_{n}^{\alpha}$ are the upper and lower endpoints of the central $(1-\alpha)\times 100\%$ prediction interval for the $n$-th data point, which are the predictive quantiles at levels $1-\frac{\alpha}{2}$ and $\frac{\alpha}{2}$ respectively; $y_{n}$ is the actual value corresponding to the $n$-th data point. The final prediction results are rewarded for having a narrower prediction interval. If the observed value falls outside the prediction interval, a penalty is incurred, and the severity of the penalty is determined by the parameter $\alpha$. Thus, a lower IS indicates better performance.
4. Methodology
To achieve effective probabilistic predictions, we propose CoST that simultaneously leverages the advantages of both deterministic and probabilistic models. Our approach involves two stages. In the first stage, the deterministic model is pretrained to predict the conditional mean that captures the primary patterns. In the second stage, the parameters of the deterministic model are frozen, and the scale-aware diffusion model, constrained by a customized fluctuation scale, is jointly trained to model residual distributions that reflect random fluctuations. Figure 2 illustrates an overview of our model.
4.1. Mean-Residual Decomposition
For urban spatiotemporal probabilistic prediction, current approaches typically employ a single probabilistic model to capture the full distribution of data, incorporating both the primary spatiotemporal patterns and the random fluctuations. However, it is challenging to model both of these distributions. Inspired by (Mardani et al., 2023) and the Reynolds decomposition in fluid dynamics (Pope, 2001), we propose to decompose the target data $\mathbf{x}^{ta}$ as follows:
(6) $\mathbf{x}^{ta}=\underbrace{\mathbb{E}[\mathbf{x}^{ta}|\mathbf{x}^{co}]}_{\begin{subarray}{c}:=\bm{\mu}(Deterministic)\end{subarray}}+\underbrace{(\mathbf{x}^{ta}-\mathbb{E}[\mathbf{x}^{ta}|\mathbf{x}^{co}])}_{\begin{subarray}{c}:=\mathbf{r}(Probabilistic)\end{subarray}},$
where $\bm{\mu}$ is the conditional mean representing the primary patterns, and $\mathbf{r}$ is the residual representing the random variations. Our core idea is that if a deterministic model can accurately predict the conditional mean, that is, $\bm{\mu}\approx\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}]$, then the probabilistic model only needs to focus on learning the simpler residual distribution, thus combining the strengths of both models to enhance the probabilistic prediction capability.
4.2. Mean Prediction via Deterministic Model
We require a deterministic model that can accurately predict the conditional mean to align with our research hypothesis, and also maintain high predictive efficiency to avoid additional increases in training and inference time. Therefore, we select the MLP-based STID model as our mean prediction module. In the first stage of training, we pretrain the model for 50 epochs to effectively capture the primary spatiotemporal patterns. Specifically, we input historical conditional data $\mathbf{x}^{co}$ into the STID model to obtain the conditional mean output $\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]$.
The STID model is pretrained by optimizing the following loss function:
(7) $\mathcal{L}_{2}=\left\|\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}]-\mathbf{x}^{ta}\right\|_{2}^{2}.$
4.3. Residual Learning via Diffusion Model
Diffusion models have achieved significant success in probabilistic modeling. In this work, we employ a diffusion model for probabilistic prediction, training it to learn the residual distribution:
(8) $\mathbf{r}_{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
Consequently, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs. (8), (2), and (3) is replaced by $\mathbf{r}_{ta}$. The residual distribution of urban spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$. Instead, it often exhibits complex spatiotemporal dependence and heterogeneity. So we consider both temporal residual learning and spatial residual learning.
Temporal Residual Learning.
For urban spatiotemporal data, the residual distribution varies at different time points. For example, fluctuations are lower at night and higher during the day, with uncertainty varying between weekends and weekdays. To model this, we incorporate the timestamp information as the condition for the denoising process. Meanwhile, the residual distribution can also be affected by sudden weather changes or public events. To capture these real-time changes, we concatenate the context data $\mathbf{x}^{co}_{0}$ with noised target residual $\mathbf{r}^{ta}_{n}$ as the input. The noise is not added to $\mathbf{x}^{co}_{0}$ and $\mathbf{r}^{ta}_{n}$ during the diffusion training and inference.
Spatial Residual Learning.
In areas with frequent traffic accidents, fluctuations tend to be more pronounced and may induce anomalous variations in adjacent regions, thus affecting their distributions. For spatial dependence modeling, the model learns a spatial embedding for each location, following the STID approach. Additionally, we propose a scale-aware diffusion process to further distinguish the heterogeneity for different regions. In this section, we detail the calculation of $\mathbf{Q}$ and how it is integrated into the scale-aware diffusion process.
(i) Customized Fluctuation Scale. Specifically, we apply the Fast Fourier Transform (FFT) to spatiotemporal sequences in the training set to quantify fluctuation levels in different regions and use the custom scale $\mathbf{Q}$ as input to account for spatial differences in residual. Specifically, we first employ FFT to extract the fluctuation components for each region within the training set. The detailed steps are as follows:
(9) $\mathbf{A}_{\mathrm{k}}=\left|\text{FFT}(\mathbf{x})_{\mathrm{k}}\right|,\quad\mathbf{{\phi}}_{\mathrm{k}}=\mathbf{\phi}\left(\text{FFT}(\mathbf{x})_{\mathrm{k}}\right),$
$\mathbf{A}_{\text{max}}=\max_{\mathrm{k}\in\left\{1,\cdots,\left\lfloor\frac{\mathbf{L}}{2}\right\rfloor+1\right\}}\mathbf{A}_{\mathrm{k}},$
$\mathcal{K}=\left\{\mathrm{k}\in\left\{1,\cdots,\left\lfloor\frac{{L}}{2}\right\rfloor+1\right\}:\mathbf{A}_{\mathrm{k}}<0.1\times\mathbf{A}_{\text{max}}\right\},$
$\mathbf{x}_{\mathbf{r}}[i]=\sum_{\mathrm{k}\in\mathcal{K}}\mathbf{A}_{\mathrm{k}}\Big{[}\cos\left(2\pi\mathbf{f}_{\mathrm{k}}i+\mathbf{\phi}_{\mathrm{k}}\right)$
$\qquad\qquad+\cos\left(2\pi\bar{\mathbf{f}}_{\mathrm{k}}i+\bar{\mathbf{\phi}}_{\mathrm{k}}\right)\Big{]},$
where $\mathbf{A}_{\mathrm{k}},\mathbf{\phi}_{\mathrm{k}}$ reprent the amplitude and phase of the $\mathrm{k}-$th frequency component. $L$ is the temporal length of the training set. $\mathbf{A}_{\text{max}}$ is the maximum amplitude among the components, obtained using the $\max$ operator. $\mathcal{K}$ represents the set of indices for the selected residual components. $\mathbf{f}_{\mathrm{k}}$ is the frequency of the $\mathrm{k}$-th component. $\bar{\mathbf{f}}_{\mathrm{k}},\bar{\mathbf{\phi}}_{\mathrm{k}}$ represent the conjugate components. $\mathbf{x}_{\mathbf{r}}$ ref to the extracted residual component of the training set. We then compute the variance $\sigma^{2}_{k}$ of the residual sequence for each location $k$ and expand it to match the shape as $\mathbf{r}^{ta}_{0}\in\mathbb{R}^{B\times K\times P}$ , where $B$ represents the batch size. And we can get the variance tensor $\mathcal{M}$:
(10) $\mathcal{M}_{b,k,p}=\sigma_{k}^{2},$
$\forall b\in\{1,\cdots,B\},\forall k\in\{1,\cdots,K\},\forall p\in\{1,\cdots,P\}.$
The residual fluctuations are bidirectional, encompassing both positive and negative variations, so we generate a random sign tensor $\mathbf{S}\in\mathbb{R}^{B\times K\times P}$ for $\mathcal{M}$, where each element $S_{b,k,p}$ of $\mathbf{S}$ is sampled from a Bernoulli distribution with $p=0.5$. The customized fluctuation scale $\mathbf{Q}$ is then defined as:
(11) $\mathbf{Q}_{b,k,p}=S_{b,k,p}\times\mathcal{M}_{b,k,p},$
$\forall b\in\{1,\cdots,B\},\forall k\in\{1,\cdots,K\},\forall p\in\{1,\cdots,P\}.$
Then $\mathbf{Q}$ is used as the input of the denoising network.
(ii) Scale-aware Diffusion Process.
The vanilla diffusion models typically model all regions as the same $\mathcal{N}(0,I)$ distribution at the end of the diffusion process, failing to distinguish the differences among regions. To further model the differences of residual distribution across various regions, we adopt the technique proposed by (Han et al., 2022) to make the residual learning region-specific conditioned on ${\mathbf{Q}}$. Specially, we have calculated the customized fluctuation scale ${\mathbf{Q}}$, and We redefined the noise distribution at the endpoint of the diffusion process as follows:
(12) $p(\mathbf{r}^{ta}_{N})=\mathcal{N}({\mathbf{Q}},I),$
Accordingly, the Eq 13 in the forward process is rewritten as:
(13) $\mathbf{r}_{n}^{ta}=\sqrt{\bar{\alpha}_{n}}\mathbf{r}_{0}^{ta}+(1-\sqrt{\bar{\alpha}_{n}})\mathbf{Q}+\sqrt{1-\bar{\alpha}_{n}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,I).$
And in the denoising process, we sample $\mathbf{r}_{N}^{ta}$ from $\mathcal{N}({\mathbf{Q}},I)$, and denoise it use Eq (2), the computation of $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ in Eq 2 is modified as:
(14) $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})=\frac{1}{\sqrt{\bar{\alpha}_{n}}}\left(\mathbf{r}_{n}^{ta}-\frac{\beta_{n}}{\sqrt{1-\bar{\alpha}_{n}}}\mathbf{\epsilon}_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})\right)+(1-\frac{1}{\sqrt{\bar{\alpha}_{n}}})\mathbf{Q}.$
This approach enables the diffusion process to be governed by the $\mathbf{Q}$ values at each region, leading to more effective utilization of the customized scale conditions.
4.4. Training and Inference
Algorithm 1 Training
1:Stage 1: Pretraining of Deterministic Model $\mathbb{E}_{\theta}$
2:repeat
3:     Estimate the conditional mean $\mathbb{E}_{\theta}[\mathbf{x}^{ta}_{0}|\mathbf{x}^{co}_{0}]$.
4:     Update $\mathbb{E}_{\theta}$ using the following loss function:
$\mathcal{L}_{2}=\left\lVert\mathbb{E}_{\theta}[\mathbf{x}^{ta}_{0}|\mathbf{x}^{co}_{0}]-\mathbf{x}^{ta}_{0}\right\rVert_{2}^{2}$
5:until The model has converged.
6:Stage 2: Training of Diffusion Model $\epsilon_{\theta}$
7:repeat
8:     Initialize $n\sim\text{Uniform}(1,\dots,N)$ and $\epsilon\sim\mathcal{N}(0,I)$.
9:     Calculate the target $\mathbf{r}_{0}^{ta}=\mathbf{x}_{0}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}_{0}|\mathbf{x}^{co}_{0}]$.
10:     Calculate noisy targets $\mathbf{r}_{n}^{ta}$ using Eq. (13).
11:     Update $\epsilon_{\theta}$ using the following loss function:
$\mathcal{L}(\theta)=\left\lVert\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})\right\rVert_{2}^{2}$
12:until The model has converged.
Algorithm 2 Inference
1:Input: Context data $\mathbf{x}_{0}^{co}$, customized fluctuation scale $\mathbf{Q}$, trained diffusion model $\epsilon_{\theta}$, trained deterministic model $\mathbb{E}_{\theta}$
2:Output: Target data $\mathbf{x}_{0}^{ta}$
3:Estimate the conditional mean $\mathbb{E}_{\theta}[\mathbf{x}^{ta}_{0}|\mathbf{x}^{co}_{0}]$
4:Sample $\mathbf{r}_{N}^{ta}$ from $\epsilon\sim\mathcal{N}(\mathbf{S},I)$
5:for $n=N$ to $1$ do
6:     Estimate the noise $\mathbf{\epsilon}_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$
7:     Calculate the $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ using Eq. (14)
8:     Sample $\mathbf{r}_{n-1}^{ta}$ using Eq. (2)
9:end for
10:Return: $\mathbf{x}_{0}^{ta}=\mathbb{E}_{\theta}[\mathbf{x}^{ta}_{0}|\mathbf{x}^{co}_{0}]+\mathbf{r}_{0}^{ta}$
Training
Our training process is a two-stage procedure. We first pretrain the deterministic model STID to enable it to predict the conditional mean. Subsequently, we train the diffusion mode to learn the distribution of residuals, where the residuals are calculated as the difference between the true values and the conditional mean predicted by the pretrained STID model with frozen parameters. The detailed training procedure is presented in Algorithm 1.
Inference
The inference process of the model consists of two paths: one utilizes the pretrained STID model to predict the conditional mean, while the other uses the diffusion model to predict the residuals. The final sample is obtained by summing the results of both paths. This process is detailed in Algorithm 2.
5. Performance Evaluations
Table 1. Short-term prediction results on five datasets in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best.
Model CellularSH BikeDC TaxiBJ CrowdBJ Los-Speed
CRPS QICE IS CRPS QICE IS CRPS QICE IS CRPS QICE IS CRPS QICE IS
D3VAE 0.856 0.105 1.73 0.785 0.157 8.77 0.433 0.160 985.7 0.668 0.099 53.6 0.119 0.089 90.5
DiffSTG 0.303 0.078 0.526 0.692 0.157 8.08 0.299 0.074 416.5 0.436 0.089 32.1 0.078 0.045 50.9
TimeGrad 0.489 0.143 0.759 0.469 0.130 5.65 0.170 0.102 213.2 0.385 0.113 48.6 0.031 0.098 20.8
CSDI 0.200 0.052 0.295 0.529 0.057 4.79 0.122 0.048 121.8 0.306 0.028 16.4 0.059 0.026 30.8
NPDiff 0.201 0.106 0.627 0.442 0.066 7.11 0.222 0.112 474.1 0.287 0.120 34.5 0.057 0.023 30.5
DyDiffusion1 0.230 0.096 0.573 0.573 0.079 6.46 0.084 0.054 99.5 - - - - - -
CoST 0.149 0.014 0.216 0.534 0.042 4.474 0.100 0.023 95.3 0.215 0.014 11.5 0.056 0.023 32.0
  • 1
    DyDiffusion can only be applied to grid-format data.
Table 2. Short-term prediction results on four datasets in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best.
Model CellularSH BikeDC TaxiBJ CrowdBJ
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
D3VAE 0.186 0.373 0.871 3.59 49.3 84.8 5.16 10.1
DiffSTG 0.066 0.103 0.770 4.02 41.8 69.4 3.68 6.63
TimeGrad 0.047 0.053 0.843 1.07 29.1 34.1 4.37 5.43
CSDI 0.044 0.075 0.592 3.10 18.2 31.6 2.71 5.51
NPDiff 0.037 0.057 0.435 1.90 26.7 52.2 2.05 3.27
DyDiffusion 0.050 0.072 0.480 1.37 12.3 18.0 - -
CoST 0.034 0.052 0.597 2.02 15.1 25.6 1.92 3.04
5.1. Experimental Settings
Datasets.
To validate the effectiveness and generalizability of our method, we use seven real-world datasets across four categories, each with different spatiotemporal granularities, covering information on crowd flow (CrowdBJ and CrowdBM), cellular network traffic (CellularNJ and CellularSH), vehicle flow (TaxiBJ and BikeDC), and traffic speed (Los-Speed). All datasets are split into training, validation, and test sets in a 6:2:2 ratio, and all datasets are standardized during training. Detailed information on the datasets can be found in Table 6.
Baselines.
we select six of the most widely recognized and state-of-the-art models in the spatiotemporal domain as our baselines, including:
  • D3VAE (Li et al., 2022): Aims at short-period and noisy time series forecasting. It combines generative modeling with a bidirectional variational auto-encoder, integrating diffusion, denoising, and disentanglement.
  • DiffSTG (Wen et al., 2023): First applies diffusion models to spatiotemporal graph forecasting. By combining STGNNs and diffusion models, it reduces prediction errors and improves uncertainty modeling.
  • TimeGrad (Rasul et al., 2021): An autoregressive model based on diffusion models. It conducts probabilistic forecasting for multivariate time series and performs well on real-world datasets.
  • CSDI (Tashiro et al., 2021):Utilizes score-based diffusion models for time series imputation. It can leverage the correlations of observed values and also shows remarkable results on prediction tasks.
  • Dyffusion (Rühling Cachay et al., 2023): A training method for diffusion models in probabilistic spatiotemporal forecasting. It combines data temporal dynamics with diffusion steps and performs well in complex dynamics forecasting.
  • NPdiff (Sheng et al., 2025): A general noise prior framework for cellular traffic prediction. It uses the data dynamics to calculate noise prior for the denoising process and achieve effective performance.
Since the dataset contains some data with a grid structure, we construct adjacency matrixes based on the adjacency relationships to support baseline models that rely on graph structures.
Metrics.
To evaluate the performance, we employed two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE, IS. For QICE, we set the number of QIs, denoted as $M_{\text{QIs}}$, to 10. For IS, we choose a confidence level of 90%, that is, $\alpha=0.1$.
Experimental Configuration.
In our experiment, for our model, we set the training maximum epoch for both the STID deterministic model and the diffusion model to 50, with early stopping based on patience of 5 for both models. During training, we set the initial learning rate to 0.001, and after 20 epochs, we adjust it to 4e-4. We use the Adam optimizer with a weight decay of 1e-6. For the diffusion model, we set the validation set sampling number to 3, and the average metric computed over these samples is used as the criterion for early stopping. For the baseline models, we set the maximum training epoch to 100 and the early stopping patience also to 5. We set the number of samples to 50 for computing the experimental results presented in the paper. In terms of model architecture, we set the encoder layer number for the STID deterministic model to 4, with an embedding dimension of 32. In the denoising network of the diffusion model, we set the encoder layer number to 8, and the embedding dimension is 128. We set the maximum diffusion steps N for the diffusion model to 50, with a linear noise schedule from $\beta_{1}=0.0001$ to the maximum $\beta_{N}=0.5$.
5.2. Short-Term Prediction
Setups.
We define the short-term prediction task as a 12-step ahead prediction based on the previous 12 steps, following (Sheng et al., 2025; Wen et al., 2023). Since the temporal granularity varies across datasets, the actual time duration corresponding to these 12 steps differs accordingly.
Results of Probabilistic Metrics.
Table 1 presents the results of probabilistic metrics for selected datasets. Due to space constraints, the remaining results can be found in the Appendix Table 7. As shown in Table 1, our model achieves the best performance across various types of datasets, including crowd flow, cellular network traffic, bike flow, taxi flow, and traffic speed. Compared to the best-performing baseline methods on each dataset, our model demonstrates an average improvement of 38% in CRPS and QICE metrics, indicating its superior ability to accurately capture the true distribution characteristics. Moreover, our model achieves a 11% improvement in the IS metric, suggesting that its prediction intervals not only maintain compactness but also exhibit higher coverage, thereby better reflecting the uncertainty of data. Although certain individual metrics may not reach the optimal level on specific datasets, our model consistently maintains performance comparable to the best methods. TimeGrad performs better on the Los-Speed dataset, likely due to its lower dynamic nature and smaller fluctuations, which suits autoregressive models better.
Table 3. Long-term prediction results on four datasets in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best.
Model CellularNJ CrowdBJ CrowdBM Los-Speed
CRPS QICE IS CRPS QICE IS CRPS QICE IS CRPS QICE IS
D3VAE 0.856 0.105 1.729 0.710 0.109 63.9 0.674 0.108 152.3 0.138 0.101 113.2
DiffSTG 0.374 0.107 0.923 0.370 0.094 31.3 0.400 0.073 67.1 0.124 0.080 104.6
TimeGrad 0.245 0.075 0.408 0.371 0.073 32.4 0.237 0.049 33.9 0.192 0.081 98.8
CSDI 0.158 0.045 0.216 0.229 0.038 12.0 0.235 0.052 33.7 0.134 0.090 59.2
NPDiff 0.204 0.102 0.611 0.288 0.114 33.6 0.331 0.111 90.8 1.366 0.126 950.4
DyDiffusion 0.308 0.086 0.550 - - - - - - - - -
CoST 0.158 0.016 0.218 0.217 0.011 11.5 0.235 0.009 31.2 0.090 0.038 65.3
Table 4. Long-term prediction results on four datasets in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best.
Model CellularNJ CrowdBJ CrowdBM Los-Speed
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
D3VAE 0.186 0.373 5.63 11.4 12.4 28.2 9.43 13.3
DiffSTG 0.078 0.125 3.04 6.37 7.59 18.8 7.77 14.2
TimeGrad 0.058 0.072 3.48 4.83 5.25 7.40 18.2 22.3
CSDI 0.035 0.057 1.99 3.64 4.64 12.4 11.3 15.0
NPDiff 0.037 0.057 2.06 3.28 5.44 13.8
DyDiffusion 0.047 0.066 - - - - - -
CoST 0.035 0.053 1.92 3.05 4.74 11.2 5.98 10.8
Results of Deterministic Metrics.
We also conduct a comparative analysis of the deterministic error between the predicted sample mean and the observation values to assess the offset between the predicted center and the ground truth. Table  2 presents a subset of the results, with more detailed information available in Appendix Table 7. The experimental results demonstrate that our method consistently outperforms others across most datasets, achieving an average reduction of 7% in MAE and 4.5% in RMSE. This indicates that, with the support of the current advanced conditional mean prediction model, our approach can effectively capture the primary patterns compared to other probabilistic baseline models.
5.3. Long-Term Prediction
Setups.
In addition to the 12-step prediction task which is the most typical setup for urban spatiotemporal prediction, we also conduct long-term prediction experiments on four. We define the long-term prediction task as a 64-step ahead prediction based on the previous 64 steps, following (Yuan et al., 2024a; Jin et al., 2023b; Nie et al., 2022).
Results of Probabilistic Metrics.
As shown in Table  3, our model demonstrates exceptional long-term prediction capability. It is worth noting that our model is based on an MLP architecture. Although some individual metrics are slightly inferior to CSDI based on Transformer, which excels at modeling long-term dependencies, our model outperforms them overall. Furthermore, in terms of training efficiency and inference speed, it significantly outperforms it, as detailed in Section 5.5.
Results of Deterministic Error Metrics.
As shown in Table 4, our model also significantly outperforms other models on deterministic metrics. Compared to the best baseline, our approach achieves an average reduction of 6.5% in MAE and 3.45% in RMSE.
Figure 3. Ablation study on the CrowdBJ and CrowdBMcomparing variants in terms of (a) CRPS and (b) IS.
5.4. Ablation Study
we conduct an ablation on each proposed module in our model to further evaluate the influence. Specifically, we constructed three variant models by progressively removing key components from the original architecture:(w/o c): The scale-aware diffusion process module is removed; (w/o q): The customized fluctuation scale is excluded as a prior condition;(w/o m): The conditional mean prediction module is eliminated, allowing the model to rely solely on the diffusion model for training. We conducted experiments on two datasets and visualize the results in terms of CRPS and IS metrics, as illustrated in Figure 3. The results demonstrate that incorporating the conditional mean prediction module significantly enhances model performance, indicating that it effectively captures the primary spatiotemporal pattern and alleviates the burden on the diffusion model to capture the full data distribution. Furthermore, incorporating the customized fluctuation Scale as a prior condition further improves prediction performance, suggesting that it provides valuable fluctuation information across different regions. Finally, the scale-aware diffusion process enables the model to better utilize this condition.
5.5. Model Efficiency
Table 5. Comparation of training time and sampling time on CellularSH dataset.
Model train time all Inference time
D3VAE 3min 27s 2min 15s
DiffSTG 24min 16s 18min 38s
TimeGrad 5min 2min
CSDI 48min 40s 38min 49s
DyDiffusion 33h 3h
CoST 2min 50s
We evaluated the training time of our model and all baseline models on the CellularSH dataset, as well as the time required to perform 50 sampling iterations on the test set. Notably, the pretraining time of the mean prediction module in our model is also included in the evaluation. The experimental results are summarized in Table 5. As shown in Table 5, our model demonstrates a significant advantage over existing probabilistic models in terms of training and inference efficiency. This efficiency is particularly valuable for real-world deployment scenarios. It is worth noting that while the CSDI model benefits from the Transformer architecture’s capability to capture spatiotemporal dependencies, its training and sampling time is considerably higher than that of other models, posing challenges for time-sensitive tasks.
5.6. Analysis of Distribution Alignment
For regions with complex distributions in the CellularSH dataset, we validate the model’s ability to capture multi-modal characteristics by comparing the predicted and true distributions using Kernel Density Estimation (KDE). As shown in Figure 4, the real distribution exhibits significant spatiotemporal multi-modality: Figure 4(a) shows three peaks, which likely represent values at different time points and values under different states at the same time points. CoST successfully captures the three peaks, while CSDI only fits two, demonstrating superior capability in modeling multimodal distributions. In Figure 4(b), the difference between the two peaks reflects strong temporal dynamics in this region. Although both models predict two peaks, the peak spacing of CoST is more consistent with the true distribution. These advantages stem from the model’s design: the diffusion model focuses on residuals, effectively capturing multimodal distribution features, while the deterministic model captures long-term trends and enhances the model’s capacity to model primary dynamic patterns.
Figure 4. KDE plots of CellularSH dataset for different regions: (a) Location 182, (b) Location 520.
Figure 5. PIT analysis on the CellularSH dataset: (a) PIT histogram and (b) PIT empirical CDF.
Figure 6. Visualizations of predictive uncertainty for both CSDI and our model on the CrowdBJ dataset. The shaded regions represent the 90% confidence interval, derived from 50 independent sampling runs. The dashed lines denote the median of the predicted values for each model.
Figure 7. PICP comparison between our model and CSDI on CrowdBM and CellularSH.
Additionally, we present the PIT (Probability Integral Transform) histogram in Figure 5 (a) and the PIT empirical cumulative distribution function (CDF) in Figure 5 (b) to visually reflect the alignment of the full distribution. Ideally, the true values’ quantiles in the predictive distribution should follow a uniform distribution, corresponding to the dashed line in Figure 5 (a). In the case of perfect calibration, the PIT CDF should closely resemble the yellow diagonal line. Clearly, our model outperforms CSDI.
5.7. Analysis of Prediction Quality
To intuitively demonstrate the effectiveness of our predictions, we visualize the prediction results on the CrowdBJ dataset, as shown in Figure 6. Here, we compare our model with the best baseline model, CSDI. On one hand, thanks to the deterministic model, our model is able to more accurately capture regular spatiotemporal patterns, as demonstrated in Figures 6 (a, c, and f). On the other hand, the diffusion model, by focusing on modeling the residual distribution, better reflects the uncertainty in predictions across different samples, as illustrated in Figures 6 (b, d, g, and h).
In addition to sample-specific analysis, we also perform a comparative analysis of the prediction intervals against the best baseline model, CSDI. We visualize the dynamic quantile errors of the CrowdBM and CellularSH prediction results in Figure 7. For each confidence level $\alpha$, we compute the corresponding quantile intervals (e.g., $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$) and calculate the proportion of real values falling within this interval, known as the Prediction Interval Coverage Probability (PICP). The closer the curve is to the diagonal (black dashed line), the better the calibration of the prediction intervals. The results show that our model achieves superior prediction interval calibration.
6. Conclusion
In this work, we propose to collaborate deterministic models and probabilistic models for effective probabilistic urban spatiotemporal predictions. By decomposing spatiotemporal data into conditional mean predicted by deterministic models and residual learned by diffusion models, bridging the gap between deterministic and probabilistic predictions, our CoST achieves dual advantages: precise capture of primary patterns and non-linear variations. Experiments across seven real-world datasets demonstrate 20% performance gains over state-of-the-art baselines. This study provides an efficient solution for urban spatiotemporal prediction that balances pattern learning and uncertainty modeling, holding significant theoretical and practical value for smart city decision support.
Appendix
Table 6. The basic information of grid-based spatio-temporal data.
Dataset City Type Temporal Period Spatial partition Interval Mean Std
TaxiBJ Beijing Taxi flow 2014/03/01 - 2014/06/30 $32\times 32$ Half an hour 111.5 139.3
BikeDC Washington, D.C. Bike flow 2010/09/20 - 2010/10/20 $20\times 20$ Half an hour 0.924 4.88
CellularSH Shanghai Cellular traffic 2014/08/01 - 2014/08/21 $32\times 28$ One hour 0.175 0.212
CellularNJ Nanjing Cellular traffic 2021/02/02 - 2021/02/22 $20\times 28$ One hour 0.842 1.30
CrowdBJ Beijing Crowd flow 2018/01/01 - 2018/01/31 $1010$ One hour 7.07 11.1
CrowdBM Baltimore Crowd flow 2019/01/01 - 2019/05/31 $403$ One hour 14.4 29.3
Los-Speed Los Angeles Traffic speed 2012/03/01 - 2012/03/07 $207$ Five minutes 59.0 12.5
Table 7. Short-term prediction results on two additional datasets in terms of both deterministic and probabilistic metrics. Bold indicates the best performance, while underlining denotes the second-best.
Model CellularNJ CrowdBM
MAE RMSE CRPS QICE IS MAE RMSE CRPS QICE IS
D3VAE 0.580 1.135 0.565 0.096 6.03 11.0 24.7 0.593 0.110 136.4
DiffSTG 0.317 0.649 0.291 0.071 3.11 8.88 21.3 0.453 0.047 68.5
TimeGrad 0.340 0.357 0.432 0.162 5.87 10.1 12.4 0.240 0.085 46.9
CSDI 0.129 0.237 0.111 0.039 0.80 7.31 19.3 0.390 0.054 61.1
NPDiff 0.123 0.175 0.128 0.133 2.22 5.42 13.7 0.331 0.119 91.2
DyDiffusion 0.222 0.357 0.196 0.080 1.80 - - - - -
CoST 0.102 0.172 0.090 0.037 0.682 5.04 12.1 0.256 0.027 37.8

### v3

@@
-Collaborative Deterministic-Diffusion Model for Probabilistic Urban Spatiotemporal Prediction
+Collaborative Deterministic-Diffusion Model for Probabilistic Spatiotemporal Prediction
 Abstract.
@@
 Existing deep learning-based deterministic forecasting models include MLP-based (Shao et al., 2022; Qin et al., 2023; Zhang et al., 2023), CNN-based (Li et al., 2017; Liu et al., 2018; Zhang et al., 2017), and RNN-based (Bai et al., 2019b; Lin et al., 2020; Wang et al., 2017, 2018) architectures, which are commonly used for their computational efficiency.
-Additionally, GNN-based models (Bai et al., 2019a, a, 2020; Geng et al., 2019; Jin et al., 2023a) have been developed to capture spatial dependencies in graph-structured data, while Transformer-based models (Chen et al., 2022, 2021a; Jiang et al., 2023; Yu et al., 2020) excel at capturing complex dynamic patterns.
+Additionally, GNN-based models (Bai et al., 2019a, 2020; Geng et al., 2019; Jin et al., 2023a) have been developed to capture spatial dependencies in graph-structured data, while Transformer-based models (Chen et al., 2022, 2021a; Jiang et al., 2023; Yu et al., 2020) excel at capturing complex dynamic patterns.
 2.3. Probabilistic Prediction
@@
 (4) $\text{QICE}:=\frac{1}{M_{\text{QIs}}}\sum_{m=1}^{M_{\text{QIs}}}\left|r_{m}-\frac{1}{M_{\text{QIs}}}\right|,$
-$r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathbb{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$
+$r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}_{y_{n}\geq\hat{y}_{n}^{\text{low}_{m}}}\cdot\mathds{1}_{y_{n}\leq\hat{y}_{n}^{\text{high}_{m}}},$
 where $\hat{y}_{n}^{\text{low}_{m}}$ and $\hat{y}_{n}^{\text{high}_{m}}$ represents the lower and upper bound of the $m-$th quantile interval for $y_{n}$.
@@
 The computation of IS can be formulated as follows:
-(5) $IS:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{\alpha}-l_{n}^{\alpha})+\frac{2}{\alpha}(l_{n}^{\alpha}-y_{n})\mathbb{1}_{y_{n}<l_{n}^{\alpha}}+\frac{2}{\alpha}(y_{n}-u_{n}^{\alpha})\mathbb{1}_{y_{n}>u_{n}^{\alpha}}\right],$
+(5) $\text{IS}:=\frac{1}{N}\sum_{n=1}^{N}\left[(u_{n}^{\alpha}-l_{n}^{\alpha})+\frac{2}{\alpha}(l_{n}^{\alpha}-y_{n})\mathds{1}_{y_{n}<l_{n}^{\alpha}}+\frac{2}{\alpha}(y_{n}-u_{n}^{\alpha})\mathds{1}_{y_{n}>u_{n}^{\alpha}}\right],$
 where $u_{n}^{\alpha}$ and $l_{n}^{\alpha}$ are the upper and lower endpoints of the central $(1-\alpha)\times 100\%$ prediction interval for the $n$-th data point, which are the predictive quantiles at levels $1-\frac{\alpha}{2}$ and $\frac{\alpha}{2}$ respectively; $y_{n}$ is the actual value corresponding to the $n$-th data point.
@@
 Diffusion models have achieved significant success in probabilistic modeling. In this work, we employ a diffusion model for probabilistic prediction, training it to learn the residual distribution:
-(8) $\mathbf{r}_{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
+(8) $\mathbf{r}^{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
 Consequently, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs.
-(8), (2), and (3) is replaced by $\mathbf{r}_{ta}$.
+(8), (2), and (3) is replaced by $\mathbf{r}^{ta}$.
 The residual distribution of urban spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$.
@@
 (12) $p(\mathbf{r}^{ta}_{N})=\mathcal{N}({\mathbf{Q}},I),$
-Accordingly, the Eq 13 in the forward process is rewritten as:
+Accordingly, the Eq (13) in the forward process is rewritten as:
 (13) $\mathbf{r}_{n}^{ta}=\sqrt{\bar{\alpha}_{n}}\mathbf{r}_{0}^{ta}+(1-\sqrt{\bar{\alpha}_{n}})\mathbf{Q}+\sqrt{1-\bar{\alpha}_{n}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,I).$
-And in the denoising process, we sample $\mathbf{r}_{N}^{ta}$ from $\mathcal{N}({\mathbf{Q}},I)$, and denoise it use Eq (2), the computation of $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ in Eq 2 is modified as:
+And in the denoising process, we sample $\mathbf{r}_{N}^{ta}$ from $\mathcal{N}({\mathbf{Q}},I)$, and denoise it use Eq (2), the computation of $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})$ in Eq (2) is modified as:
 (14) $\mu_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})=\frac{1}{\sqrt{\bar{\alpha}_{n}}}\left(\mathbf{r}_{n}^{ta}-\frac{\beta_{n}}{\sqrt{1-\bar{\alpha}_{n}}}\mathbf{\epsilon}_{\theta}(\mathbf{r}_{n}^{ta},n|\mathbf{x}_{0}^{co})\right)+(1-\frac{1}{\sqrt{\bar{\alpha}_{n}}})\mathbf{Q}.$
@@
 NPDiff 0.201 0.106 0.627 0.442 0.066 7.11 0.222 0.112 474.1 0.287 0.120 34.5 0.057 0.023 30.5
-DyDiffusion1 0.230 0.096 0.573 0.573 0.079 6.46 0.084 0.054 99.5 - - - - - -
+DYffusion1 0.230 0.096 0.573 0.573 0.079 6.46 0.084 0.054 99.5 - - - - - -
 CoST 0.149 0.014 0.216 0.534 0.042 4.474 0.100 0.023 95.3 0.215 0.014 11.5 0.056 0.023 32.0
   • 1
-    DyDiffusion can only be applied to grid-format data.
+    DYffusion can only be applied to grid-format data.
 Table 2. Short-term prediction results on four datasets in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best.
@@
 NPDiff 0.037 0.057 0.435 1.90 26.7 52.2 2.05 3.27
-DyDiffusion 0.050 0.072 0.480 1.37 12.3 18.0 - -
-CoST 0.034 0.052 0.597 2.02 15.1 25.6 1.92 3.04
+DYffusion 0.050 0.072 0.480 1.37 12.3 18.0 - -
+CoST 0.034 0.052 0.492 1.76 15.1 25.6 1.92 3.04
 5.1. Experimental Settings
@@
 Metrics.
-To evaluate the performance, we employed two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE, IS.
+To evaluate the performance, we employed two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE and IS.
 For QICE, we set the number of QIs, denoted as $M_{\text{QIs}}$, to 10.
@@
 NPDiff 0.204 0.102 0.611 0.288 0.114 33.6 0.331 0.111 90.8 1.366 0.126 950.4
-DyDiffusion 0.308 0.086 0.550 - - - - - - - - -
+DYffusion 0.308 0.086 0.550 - - - - - - - - -
 CoST 0.158 0.016 0.218 0.217 0.011 11.5 0.235 0.009 31.2 0.090 0.038 65.3
@@
 CSDI 0.035 0.057 1.99 3.64 4.64 12.4 11.3 15.0
-NPDiff 0.037 0.057 2.06 3.28 5.44 13.8
-DyDiffusion 0.047 0.066 - - - - - -
+NPDiff 0.037 0.057 2.06 3.28 5.44 13.8 46.0 58.3
+DYffusion 0.047 0.066 - - - - - -
 CoST 0.035 0.053 1.92 3.05 4.74 11.2 5.98 10.8
@@
 Furthermore, in terms of training efficiency and inference speed, it significantly outperforms it, as detailed in Section 5.5.
-Results of Deterministic Error Metrics.
+Results of Deterministic Metrics.
 As shown in Table 4, our model also significantly outperforms other models on deterministic metrics.
@@
 CSDI 48min 40s 38min 49s
-DyDiffusion 33h 3h
+DYffusion 33h 3h
 CoST 2min 50s
-We evaluated the training time of our model and all baseline models on the CellularSH dataset, as well as the time required to perform 50 sampling iterations on the test set.
+We evaluated the training and time of our model and all baseline models on the CellularSH dataset, as well as the time required to perform 50 sampling iterations on the test set.
 Notably, the pretraining time of the mean prediction module in our model is also included in the evaluation.
@@
 As shown in Figure 4, the real distribution exhibits significant spatiotemporal multi-modality: Figure 4(a) shows three peaks, which likely represent values at different time points and values under different states at the same time points.
-CoST successfully captures the three peaks, while CSDI only fits two, demonstrating superior capability in modeling multimodal distributions.
+CoST successfully captures the three peaks, while CSDI only fits two, demonstrating superior capability in modeling multi-modal distributions.
 In Figure 4(b), the difference between the two peaks reflects strong temporal dynamics in this region.
 Although both models predict two peaks, the peak spacing of CoST is more consistent with the true distribution.
-These advantages stem from the model’s design: the diffusion model focuses on residuals, effectively capturing multimodal distribution features, while the deterministic model captures long-term trends and enhances the model’s capacity to model primary dynamic patterns.
+These advantages stem from the model’s design: the diffusion model focuses on residuals, effectively capturing multi-modal distribution features, while the deterministic model captures long-term trends and enhances the model’s capacity to model primary dynamic patterns.
 Figure 4. KDE plots of CellularSH dataset for different regions: (a) Location 182, (b) Location 520.
@@
 NPDiff 0.123 0.175 0.128 0.133 2.22 5.42 13.7 0.331 0.119 91.2
-DyDiffusion 0.222 0.357 0.196 0.080 1.80 - - - - -
+DYffusion 0.222 0.357 0.196 0.080 1.80 - - - - -
 CoST 0.102 0.172 0.090 0.037 0.682 5.04 12.1 0.256 0.027 37.8